OpenStudy (anonymous):

A cylinder and a rectangular prism have the same volume and same height. The base of the prism is a square with side length 9 cm. What is the approximate radius of the cylinder? A. 4.5 cm B. 5.1 cm C. 12.9 cm D. 25.8 cm AND PLEASE EXPLAIN BECAUSE IM SO CONFUSED

OpenStudy (anonymous):

Volume of prism is 1/3 * area of base* height in this case base is square so we will put the area of square and volume of cylinder is pi * r^2 * height equate both the volumes u will surely get your answer

OpenStudy (anonymous):

and so 9 is the area ? @niksva

OpenStudy (anonymous):

area of square = (side)^2 here side is 9 so can u determine the area?

OpenStudy (anonymous):

whats the ^ mean

OpenStudy (anonymous):

9 multiplied by two which is 81? the area

OpenStudy (anonymous):

9^2 means 9 raised to power 2 which is 81

OpenStudy (anonymous):

oh yeah 9 squared

OpenStudy (anonymous):

can you draw something showing how to find the approximate radius

OpenStudy (anonymous):

now equate the volume in this this manner \[\pi*r^2h = \frac{ 1 }{ 3 }* area of \square * h\]

OpenStudy (anonymous):

now you can determine radius

OpenStudy (anonymous):

whats the f mean im in the 6th grade @niksva

OpenStudy (anonymous):



OpenStudy (anonymous):

your equation says that i have to multiply the radius and height when i dont even know the radius and height @niksva

OpenStudy (anonymous):

@jim_thompson5910 can u help

OpenStudy (anonymous):

can anyone help

OpenStudy (anonymous):

height of prism and cylinder are equal so height will be cancelled out and u need to determine the radius

OpenStudy (anonymous):

is it B. 5.1

OpenStudy (anonymous):

oohoho its D

OpenStudy (anonymous):

i was just gonna say that

OpenStudy (anonymous):

because...

OpenStudy (anonymous):

9 * 9 is 81 then you divide 81 by 3.14

OpenStudy (anonymous):

\[22/7 * r^2 = 1/3 *81\]

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

sorry we are not going to multiply by 1/3 in prism r^2 = 81/3.14

OpenStudy (anonymous):

divide it now and then take the square root u will get the radius

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so answer will be B

OpenStudy (anonymous):

thx @niksva

OpenStudy (anonymous):

and formulae for volume of prism is area of base * height pls do make that correction

OpenStudy (anonymous):

k but its still B right @niksva

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

one more question its small

OpenStudy (anonymous):

the exact answer is closest to the answer than the approximate right

OpenStudy (anonymous):

im pretty sure it is

OpenStudy (anonymous):

exact answer is 5.07 cm approximate answer is 5.10 cm

OpenStudy (anonymous):

forget it